1) how do you calculate savings given the income that is 500, output 500 and consumption 200.
2) how do you calculate consumption given the income that is 750, output 750 and savings 400
3) how do you calculate consumption given the income that is 800, output 800 and savings 230
1
Expert's answer
2020-10-30T13:06:18-0400

1) If income is 500, output is 500, and consumption is 200, then savings are S = 500 - 200 = 300.

2) If income is 750, output is 750, and savings are 400, then consumption is C = 750 - 400 = 350.

3) If income is 800, output is 800, and savings are 230, then consumption is C = 800 - 230 = 570.